What is diffrential form of centre force.
1
Expert's answer
2020-12-04T11:19:09-0500

Answer

Differential form of Central force is given by

"\frac{d^2u}{d\theta^2}+u=-\frac{F}{mh^2u^2}\\where"

"u=\frac{1}{r}"